The formula for calculating the amount of money returned for an initial deposit into a bank account or CD (certificate of deposit) is given by
A is the amount of the return.
P is the principal amount initially deposited.
r is the annual interest rate (expressed as a decimal).
n is the number of compound periods in one year.
t is the number of years.
Carry all calculations to six decimals on each intermediate step, then round the final answer to the nearest cent.
Suppose you deposit $4,000 for 8 years at a rate of 7%.
calculate the return(A) if the bank compounds annually (n=1). round the answer to the hundreth's place

Answer by checkley71(8403)

A=P(1+R/N)^NT
A=4,000(1+.07/1)^1*8
A=4,000(1.07)^8
A=4,000*1.718186
A=$6,872.74 IS THE TOTAL AMOUNT OF THE INVESTMENT AFTER 8 YEARS.